# Combining Serverless and High-Performance Computing Paradigms to support ML Data-Intensive Applications

### v1

@@ I Introduction and Motivation
Function as a Service (FaaS) is a rapidly evolving paradigm in cloud computing applications. In FaaS, users focus on writing code divided into individual functions. This approach allows users to focus on the application code itself, rather than on deploying and managing the underlying compute and storage infrastructure. Serverless functions also provide elasticity and fine-grained billing capabilities. However, serverless currently lacks efficient communication through message passing for large-scale parallel data processing applications, such as Cylon[copik2023fmi]. To address this limitation, Network Address Translation (NAT) TCP Hole Punching offers a high-performance and cost-effective alternative to other methods, such as object or in-memory storage used on AWS Lambda to circumvent communication challenges[moyer2021punching]. Copik et al. detail performance experiments of the FMI library in the paper ”FMI: Fast and Cheap Message Passing for Serverless Functions.” Performance improvements for FMI usage are observed to be between 105 and 1024 times better than that of AWS DynamoDB. Additionally, the authors detail experiments demonstrating performance converging comparably to FMI and MPI on EC2 around 8 nodes and note improvements in performance for distributed machine learning FaaS applications 162 times and cost reductions up to 397x times[copik2023fmi].
@@
In this paper, we will explore our high-performance serverless, a runtime engine designed to optimize the execution of deep learning applications in high-performance computing (HPC) and cloud environments. We introduce a serverless communicator design that leverages NAT Hole Punching to provide a high-performance alternative to intermediate storage solutions like AWS S3. This enables function-to-function bidirectional communication, a capability not natively supported by AWS Lambda. This is also an ideal choice for AWS Fargate, serverless Docker on AWS ECS, based on the difficulties in implementing high-performance communication using libraries such as UCX, MPI, and Gloo. We introduce and incorporate Cylon to achieve an existing goal of parallel computing, making it achievable without user intervention. This is accomplished by utilizing libraries of previously parallelized operators (such as join and all the other Pandas operators, as well as deep learning). We will explain the advantages of executing workloads within Docker containers on AWS using ECS tasks and discuss how this approach can be extended to HPC environments utilizing Singularity or Apptainer so that we can execute derivative experiments.
@@ II Related Works
Another related project, Wukong, offers a similar capability, including support for AI/ML inference, but was built on serverless FaaS where AWS manages provisioning, scaling, and other undifferentiated tasks. This DAG execution framework has demonstrated near-ideal scalability by executing jobs approximately 68 times faster while achieving nearly 92 percent cost savings compared to NumPyWren. Wukong utilizes Redis as a key-value store for intermediate and final storage, addressing the data parallelism and communication performance issues described in related scientific works.
@@ III-E Serverless High-Performance BSP Architecture
Figure 6 illustrates our architecture. Our design calls for the use of an AWS Step Function to encapsulate the AWS Serverless Execution Engine, as shown in Figure 7. AWS Step Functions are a Software as a Service (SaaS) feature that facilitates the orchestration of AWS services into serverless workflows. The Lambda: init function receives a payload containing various parameters, such as the number of rows, world size, number of iterations, target S3 bucket, script, and output paths. We utilize a specific S3 bucket to store invocation results in a folder using the Boto3 library. This approach is advantageous over parsing the data output from log files in AWS CloudWatch, a SaaS metrics repository created and utilized by AWS services, as it would be cumbersome. The Lambda invoke state forwards this input JSON payload to a fmi_init AWS Lambda function. This function validates the availability of the rendezvous server. It generates an N-sized array of JSON payloads, which is then sent to the Map, ExtractAndInvokeLambda states, which is configured to use distributed processing mode to enable highly parallel execution. Next, each generated payload is forwarded to the Lambda: Invoke state from the pass ExtractAndInvokeLambda state. This state passes the input to output, executing the fmi_executor AWS Lambda function. The runtime of the fmi_executor function includes the high-performance MLSys runtime described earlier, which consists of a communicator that supports direct communication over a NAT gateway described earlier. The incoming payload specifies the location of the target script in the S3 bucket. Upon execution of the Lambda function, the contents of the specified S3 bucket are retrieved and persisted in temporary ephemeral storage. Subsequently, Python 3 is used to execute the script used for our experiments. Our design supports both a single script and a folder containing Python scripts and modules.
@@
During execution, the fmi_executor connects with the Rendezvous Server referenced in Figure 6. Once the second function in the pair connects to the Rendezvous server, both servers receive the destination function’s public address. The Python script continues to run until the collective operations are performed. Our design also includes using the cloudmesh stopwatch library to facilitate logging and benchmarking the execution time of our experiments[cloudmeshcommon].
@@ VI Conclusion
Second, the reference FMI library did not include an All-To-All collective operation that we needed for Cylon communicator integration. Second, it was necessary to include variable allgather, variable gather, support for non-blocking operations, retries for socket connection failures, a ping/ping capability to prevent eager termination of sockets, and rank calculation using atomic counters. An initial challenge was coordination between nodes and associated race conditions. Rank-based ordering for blocking operations using Redis locks was incorporated into our design to address race conditions observed during blocking socket operations. Note that FMI, in its original form, only supports non-blocking operations, and our design called for non-blocking I/O. These additions resulted in improvements and a robust communicator that can be leveraged across serverless platforms.

### v2

@@ I Introduction and Motivation
Function as a Service (FaaS) is a rapidly evolving paradigm in cloud computing applications. In FaaS, users focus on writing code divided into individual functions. This approach allows users to focus on the application code itself, rather than on deploying and managing the underlying compute and storage infrastructure. Serverless functions also provide elasticity and fine-grained billing capabilities. However, serverless currently lacks efficient communication through message passing for large-scale parallel data processing applications, such as Cylon[copik2023fmi]. To address this limitation, Network Address Translation (NAT) TCP Hole Punching offers a high-performance and cost-effective alternative to other methods, such as object or in-memory storage used on AWS Lambda to circumvent communication challenges[moyer2021punching]. Copik et al. detail performance experiments of the FMI library in the paper ”FMI: Fast and Cheap Message Passing for Serverless Functions.” Performance improvements for FMI usage are observed to be between 105 and 1024 times better than that of AWS DynamoDB. Additionally, the authors detail experiments demonstrating performance converging comparably to FMI and MPI on EC2 around 8 nodes and note improvements in performance for distributed machine learning FaaS applications 162 times and cost reductions up to 397 times[copik2023fmi].
@@
In this paper, we will explore our high-performance serverless, a runtime engine designed to optimize the execution of deep learning applications in high-performance computing (HPC) and cloud environments. We introduce a serverless communicator design that leverages NAT Hole Punching to provide a high-performance alternative to intermediate storage solutions like AWS S3. This enables function-to-function bidirectional communication, a capability not natively supported by AWS Lambda. This is also an ideal choice for AWS Fargate, serverless Docker on AWS ECS, based on the difficulties in implementing high-performance communication using libraries such as UCX, MPI, and Gloo. We introduce and incorporate Cylon to achieve an existing goal of parallel computing, making it achievable without user intervention. This is accomplished by utilizing libraries of previously parallelized operators (such as join and all the other Pandas operators, as well as deep learning). We will explain the advantages of executing workloads within Docker containers on AWS using ECS tasks and discuss how this approach can be extended to HPC environments utilizing Singularity or Apptainer so that it is possible to execute derivative experiments.
@@ II Related Works
Another related project, Wukong, offers a similar capability, including support for AI/ML inference, but was built on serverless FaaS where AWS manages provisioning, scaling, and other undifferentiated tasks. This DAG execution framework has demonstrated near-ideal scalability by executing jobs approximately 68 times faster while achieving nearly 92 percent cost savings compared to NumPyWren. Wukong utilizes Redis as a key-value store for intermediate and final storage, addressing the data parallelism and communication performance issues described in related scientific works[carver2020wukong].
@@ III-E Serverless High-Performance BSP Architecture
Figure 6 illustrates our architecture. Our design calls for the use of an AWS Step Function to encapsulate the AWS Serverless Execution Engine, as shown in Figure 7. AWS Step Functions are a Software as a Service (SaaS) feature that facilitates the orchestration of AWS services into serverless workflows. The Lambda: init function receives a payload containing various parameters, such as the number of rows, world size, number of iterations, target S3 bucket, script, and output paths. We utilize a specific S3 bucket to store invocation results in a folder using the Boto3 library. This approach is advantageous over parsing the data output from log files in AWS CloudWatch, a SaaS metrics repository created and utilized by AWS services, as it would be cumbersome. The Lambda invoke state forwards this input JSON payload to a cylon_init AWS Lambda function. This function validates the availability of the rendezvous server. It generates an N-sized array of JSON payloads, which is then sent to the Map, ExtractAndInvokeLambda states, which is configured to use distributed processing mode to enable highly parallel execution. Next, each generated payload is forwarded to the Lambda: Invoke state from the pass ExtractAndInvokeLambda state. This state passes the input to output, executing the fmi_executor AWS Lambda function. The runtime of the cylon_executor function includes the high-performance MLSys runtime described earlier, which consists of a communicator that supports direct communication over a NAT gateway described earlier. The incoming payload specifies the location of the target script in the S3 bucket. Upon execution of the Lambda function, the contents of the specified S3 bucket are retrieved and persisted in temporary ephemeral storage. Subsequently, Python is used to execute the script used for our experiments. Our design supports both a single script and a folder containing Python scripts and modules.
@@
During execution, the cylon_executor connects with the Rendezvous Server referenced in Figure 6. Once the second function in the pair connects to the Rendezvous server, both servers receive the destination function’s public address. The Python script continues to run until the collective operations are performed. Our design also includes using the cloudmesh stopwatch library to facilitate logging and benchmarking the execution time of our experiments[cloudmeshcommon].
@@ VI Conclusion
Second, the reference FMI library did not include an All-To-All collective operation that we needed for Cylon communicator integration. Second, it was necessary to include variable allgather, variable gather, support for non-blocking operations, retries for socket connection failures, a ping/ping capability to prevent eager termination of sockets, and rank calculation using atomic counters. An initial challenge was coordination between nodes and associated race conditions. Rank-based ordering for blocking operations using Redis locks was incorporated into our design to address race conditions observed during blocking socket operations. Note that FMI, in its original form, only supports blocking operations, and our design called for non-blocking I/O. These additions resulted in improvements and a robust communicator that can be leveraged across serverless platforms.